OpenStudy (anonymous):

If 14y - 4 = 24y + 26, then 12y = ?

OpenStudy (anonymous):

12y=22y+4 Because 14y-2y-4+4=24y-2y+4

OpenStudy (anonymous):

12y=22y+30

OpenStudy (anonymous):

MarinaDL is correct, I didn't see the last of the equation.

OpenStudy (anonymous):

24y-14y=-4-26 10y=-30 y=-3 12y=-36